# Empowering Source-Free Domain Adaptation via MLLM-Guided Reliability-Based Curriculum Learning

### v2

@@ Abstract
Source-Free Domain Adaptation (SFDA) aims to adapt a pre-trained source model to a target domain using only unlabeled target data. Current SFDA methods face challenges in effectively leveraging pre-trained knowledge and exploiting target domain data. Multimodal Large Language Models (MLLMs) offer remarkable capabilities in understanding visual and textual information, but their applicability to SFDA poses challenges such as instruction-following failures, intensive computational demands, and difficulties in performance measurement prior to adaptation. To alleviate these issues, we propose Reliability-based Curriculum Learning (RCL), a novel framework that integrates multiple MLLMs for knowledge exploitation via pseudo-labeling in SFDA. Our framework incorporates Reliable Knowledge Transfer, Self-correcting and MLLM-guided Knowledge Expansion, and Multi-hot Masking Refinement to progressively exploit unlabeled data in the target domain. RCL achieves state-of-the-art (SOTA) performance on multiple SFDA benchmarks, e.g., +9.4% on DomainNet, demonstrating its effectiveness in enhancing adaptability and robustness without requiring access to source data. Our code is available at https://github.com/Dong-Jie-Chen/RCL
$\ast$ Equal contribution.
Correspondence to: Dongjie Chen $<$cdjchen@ucdavis.edu$>$.
@@ 1 Introduction
(b) Accuracy vs. Model Size (Number of Parameters, log scale).
@@
Source-Free Domain Adaptation (SFDA) aims to adapt a pre-trained model to a new target domain without requiring access to labeled source data. This setting is particularly useful when privacy, storage, or proprietary constraints prevent access to source datasets. One of the key aspects that make SFDA effective is the use of externally pre-trained models. Earlier works utilized models pre-trained on ImageNet [20], while more recent works shift towards using large-scale pre-trained models like CLIP [7, 15, 33, 41]. These models capture rich, transferable representations that bridge the domain gap. However, the importance of pre-trained knowledge has been less explored in previous works [20].
Recently, multimodal large language models (MLLMs) such as LLaVA [24] and InstructBLIP [4] have set new performance records in tasks like visual question answering (VQA), detailed description, and complex reasoning retrieval [21, 25]. With their strong generalization ability, MLLMs have been widely adopted for efficient fine-tuning in downstream tasks, particularly in scenarios with limited training data. However, a critical observation is that the datasets used to train MLLMs are typically inaccessible due to proprietary restrictions or their scale and complexity. Consequently, utilizing MLLMs for downstream tasks inherently aligns with the principles of SFDA, as they transfer knowledge without access to the source training data. This necessitates the exploration of SFDA with MLLMs, especially in specialized domains such as medical and industrial applications, where source data is often sensitive or restricted, and MLLMs could have transformative impacts.
Figure 2: An overview of our proposed Reliability-based Curriculum Learning (RCL) framework.
Nevertheless, the application of MLLMs in SFDA is still limited. One of the primary challenges in employing MLLMs for SFDA is that MLLMs are not suitable for zero-shot classification tasks. As illustrated in Figure 3, MLLMs often fail to follow classification instructions as they are primarily designed for text generation. Remarkably, with improvements on instruction following (with STS as described in Sec. 3.1), these MLLMs achieve superior results without having been exposed to any images from either the source or target domain. As shown in Figure 1(a), InstructBLIP and LLaVA demonstrate better performance compared to models pre-trained on ImageNet (C-SFDA [13]) or distilling knowledge from CLIP (DIFO-C-ViT [41]). Based on this observation, we hypothesize that pre-trained knowledge can play an equally or more important role than pre-trained models in improving the performances of SFDA.
While MLLMs can inject valuable pre-trained knowledge, their applications to SFDA still face two major challenges. First, their inference process is time-consuming and computationally intensive, which prevents their wide adoption. Second, MLLMs exhibit variability in zero-shot SFDA performance, (see Figure 1(b)). Their effectiveness depends not only on model size but also on differences in their pretraining datasets and architectures, leading to diverse knowledge representations. Consequently, relying on a single MLLM for SFDA can result in suboptimal adaptation, particularly when the target domain shifted with MLLM’s training data in a large margin. Additionally, fine-tuning MLLMs for specific domains requires substantial computational resources and a GPT-4-based instruction-following dataset [4, 25], creating scalability challenges across diverse downstream tasks. This underscores the need for efficient methods that utilize multiple MLLMs for knowledge transfer while preserving scalability for domain-specific adaptations.
To overcome these issues, we first propose Semantic Textual Similarity (STS), a method that reformulates image classification as a VQA task by prompting MLLMs with predefined class names. We observe that MLLMs sometimes fail to strictly follow instructions (see Fig 3), producing open-ended responses influenced by their pretraining. STS aligns these outputs with target class labels, ensuring MLLMs function reliably for classification. As shown in Figure 1(a), zero-shot MLLMs with our STS can rival SOTA SFDA methods without any fine-tuning, making them an attractive choice for adaptation. Beyond using a single MLLM, a fundamental limitation remains: knowledge transfer from one MLLM is inherently constrained—pseudo-labels derived from a single model do not fully capture the diverse reasoning capabilities of multiple MLLMs. A naive approach would be to directly distill knowledge from an ensemble of MLLMs, but this leads to inconsistent pseudo-labels and over-reliance on a single teacher’s biases. Instead, we introduce Reliability-based Curriculum Learning (RCL), a novel multi-MLLM distillation framework that incorporates reliability-driven learning. RCL integrates multiple MLLMs for SFDA through agreement-driven knowledge distillation. Unlike traditional multi-teacher knowledge distillation (MTKD), which assumes that teacher models provide uniformly reliable guidance, RCL introduces the concept of pseudo-label reliability based on MLLM agreement. This is particularly critical because MLLMs perform zero-shot classification, making their outputs inherently uncertain.
To this end, we: (1) Quantify pseudo-label reliability by measuring agreement among multiple MLLMs, ensuring that knowledge transfer is guided by robust, consensus-driven supervision. (2) Design a structured curriculum learning strategy that progressively incorporates pseudo-labels based on their reliability: the model first learns from high-confidence samples, then integrates less reliable pseudo-labels with adaptive correction, and finally incorporates uncertain or previously unlabeled samples through Multi-hot Masking Refinement (MMR). This staged approach prevents early overfitting to unreliable labels and enables a self-correcting adaptation process. (3) Introduce MMR within the curriculum learning framework to address uncertainty by refining noisy pseudo-labels rather than discarding them, ensuring that all available target data contributes to model adaptation. This refinement process leverages multi-hot masking and consistency regularization to improve pseudo-label accuracy, allowing the model to effectively utilize ambiguous samples while mitigating errors.
Leveraging MLLM knowledge in a structured adaptation framework, RCL achieves SOTA SFDA performance on Office-Home, DomainNet, and VisDA, surpassing single-MLLM zero-shot and recent SFDA methods. Our results highlight that a multi-MLLM distillation approach, when guided by reliability and curriculum learning, enhances adaptation robustness without requiring fine-tuning or computationally expensive retraining. While our primary focus is SFDA, the principles of reliability-driven multi-teacher distillation and curriculum learning could extend beyond SFDA to broader applications in knowledge transfer, model adaptation, and robust AI training.
@@ 2 Related Work
Source-Free Domain Adaptation. SFDA adapts a pre-trained source model to a target domain using unlabeled target data, making pseudo-labeling a key technique [29, 3, 22]. Some works improve pseudo-labeling by leveraging target data structure [10, 35, 37] or aligning source and target distributions [46, 5]. Others explore synthetic target-style data generation [19, 42, 49]. Our approach leverages MLLMs’ zero-shot capabilities for pseudo-labeling and performs multi-teacher distillation through a curriculum learning framework to enhance adaptation.
VLMs/MLLMs. Pretrained vision-language models (VLMs) like CLIP [30] and ALIGN [9] capture vision-language-aligned features, while recent MLLMs (e.g., LLaVA [24], InstructBLIP [4]) enhance multimodal understanding via LLM backbones, enabling strong zero-shot capabilities. While some works use CLIP for domain adaptation, they require labeled source data [16, 50, 43] or fine-tuning [15]. DIFO [41], the latest SOTA work, adapts CLIP for target-domain learning by iteratively customizing it with prompt learning and distilling its knowledge into a task-specific model. While effective, this approach requires adapting a large VLM, whereas our method leverages multiple MLLMs in a zero-shot manner without finetuning.
Multi-Teacher Knowledge Distillation (MTKD). MTKD extends traditional KD by distilling knowledge from multiple teachers [11, 28], typically assuming labeled source data and fixed teacher ensembles [44, 18]. Wu et al. [44] proposed aligning teacher outputs to ensure consistent supervision, while Li et al. [18] introduced knowledge integration to enhance student generalization beyond simple imitation. Kang et al. [12] explored heterogeneous teacher architectures in recommendation tasks, demonstrating the benefits of aggregating diverse model types for improved performance. Ding et al. [6] studied balancing teacher ensemble size by learning a categorical distribution for stochastic teacher selection, optimizing the trade-off between capacity and efficiency. Unlike prior approaches that assume labeled source data and static teacher ensembles, our method leverages MLLMs for SFDA without source access. RCL uniquely refines pseudo-labels via curriculum learning, incorporating reliability scoring to adapt MLLM knowledge from multiple teachers—to the best of our knowledge, this has not been explored before. The proposed MMR (Sec. 4.3) enhances robustness by mitigating pseudo-label noise across domains. Without stochastic teacher selection, RCL ensures structured, self-correcting adaptation, achieving strong SFDA performance even under varying pseudo-label confidence.
Figure 3: Pseudo-labeling with MLLMs. Directly prompting MLLMs for classification can lead to failures: we propose Semantic Textual Similarity (STS) to correct pseudo-labelling.
3 Pseudo-labeling and Reliability Measurement with MLLMs
@@
3.1 Pseudo-labeling with MLLMs
We leverage multiple MLLMs for initial target-image labeling. As MLLMs are primarily designed for text generation, their responses may deviate from classification requirements. Therefore, we design prompts to repurpose MLLMs for class label prediction. Figure 3 shows how we reframe VQA as a zero-shot classification task using MLLMs. We design the prompt to incorporate all class names and a question instructing MLLMs to select the most appropriate match. This prompt and the image $x_{t}^{i}$ are then fed to multiple pre-trained MLLMs, such as LLaVA [24, 25], InstructBLIP [4], and ShareGPT4V [2]. Each MLLM generates a text output $T_{1}^{i},T_{2}^{i},\ldots,T_{M}^{i}$, where $M$ is the number of MLLMs.
However, as shown in Figure 3, MLLMs sometimes fail to follow prompts, generating responses beyond classification constraints. This happens when MLLMs rely on prior knowledge rather than selecting from provided options (e.g., predicting ‘Audi’ instead of ‘car’). Unlike MTKD, which distills structured soft labels, our approach must handle inherently diverse and inconsistent MLLM outputs. To mitigate this, we propose Semantic Textual Similarity (STS) to align outputs with class names, ensuring pseudo-label consistency.
We derive pseudo-labels by computing STS between class names and MLLM-generated text. Formally, for the $m$-th MLLM and the $i$-th image $x_{t}^{i}$, the pseudo-label $\hat{y}^{mi}$ is determined by:
$\hat{y}^{mi}=\operatorname*{argmax}_{c}\text{STS}(T_{m}^{i},T_{t}^{c}),$ (1)
where $T_{t}^{c}$ represents the name of the $c$-th class. The STS between two text sequences $T_{1}$ and $T_{2}$ is computed as:
$\text{STS}(T_{1},T_{2})=\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{\|\mathbf{v}_{1}\|_{2}\|\mathbf{v}_{2}\|_{2}}-1,$ (2)
where $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ are vector representations (using [31]) of $T_{1}$ and $T_{2}$, respectively. STS refines pseudo-labels by aligning MLLM outputs with class semantics, even when instructions are ignored. To facilitate pseudo-labeling, we design specific prompt templates for different MLLMs: (1) LLaVA models: "Question: What is the closest name from this list to describe the object in the image? Return the name only. <class names>" (2) ShareGPT4V models: "Question: What is the closest name from this list to describe the object in the image? List: <class names> Return the closest name from the list only. Use *exact* names from the list only. Answer:" (3) InstructBLIP models: "Question: What is the closest name from this list to describe the object in the image? <class names>. Use the closest name from the list only. Pick the answer from the list only. Answer:"
@@ 3.2 Consensus-based Reliability Measurement
Pseudo-labels from different MLLMs may vary for the same target sample. This disagreement raises a key question: how can we measure the reliability of the pseudo-labels from multiple MLLMs? While STS helps correct deviations from instructions, it does not assess pseudo-label reliability, as it cannot detect when MLLMs generate incorrect labels. To address this, we propose a consensus-based reliability metric for pseudo-labels.
We define a reliability score $\mathcal{R}(x_{t}^{i})$ for each target domain sample $x_{t}^{i}$ based on agreement among MLLM-assigned pseudo-labels:
@@
where $\mathbbm{1}(\cdot)$ is the indicator function. $\mathcal{R}(x_{t}^{i})$ quantifies the proportion of MLLM pairs that assign the same pseudo-label to $x_{t}^{i}$. Using $\mathcal{R}(x_{t}^{i})$, we categorize $\mathcal{D}_{t}$ into three subsets: (1) Reliable ($\mathcal{D}_{\textit{R}}$): All MLLMs agree ($\mathcal{R}(x_{t}^{i})=1$). (2) Less Reliable ($\mathcal{D}_{\textit{LR}}$): Partial agreement ($10<\mathcal{R}(x_{t}^{i})<$ 1). (3) Unreliable ($\mathcal{D}_{\textit{UR}}$): No agreement ($\mathcal{R}(x_{t}^{i})=0$). Thus, the target dataset is partitioned as: $\mathcal{D}_{t}=\{\mathcal{D}_{\textit{R}},\mathcal{D}_{\textit{LR}},\mathcal{D}_{\textit{UR}}\}.$ As shown in Figure 4, higher reliability scores correlate with improved pseudo-label accuracy, validating our consensus-based reliability metric. Furthermore, Figure 4 shows the accuracy and distribution of the pseudo-label at different levels of reliability $\mathcal{R}(x_{t}^{i})$, highlighting the disagreement between LLaVA, InstructBLIP, and ShareGPT4V. For samples with $\mathcal{R}(x_{t}^{i})>0$, accuracy is computed using a majority vote across MLLMs, while for $\mathcal{R}(x_{t}^{i})=0$, we report the accuracy of each individual MLLM.

### v3

@@ Abstract
Existing SFDA methods struggle to fully use pre-trained knowledge and often rely on a single model’s predictions or handcrafted prompts, limiting robustness under domain shift. Multimodal Large Language Models (MLLMs) offer a promising alternative: they encode rich visual-semantic knowledge and generalize well without task-specific tuning. However, their use in SFDA is hindered by instruction-following failures, inconsistent outputs, and high inference costs. We propose Reliability-based Curriculum Learning (RCL), a novel framework that distills robust supervision from multiple frozen MLLMs into a compact target model. RCL organizes adaptation as a three-stage curriculum that progressively incorporates pseudo-labels based on inter-model agreement and model confidence, enabling stable and noise-aware training. Our approach achieves state-of-the-art performance on standard SFDA datasets, Office-Home, DomainNet-126, and VisDA-C, outperforming zero-shot MLLMs, their ensembles, all without accessing source data or tuning foundation models. Code is available at https://github.com/Dong-Jie-Chen/RCL.
@@ 1 Introduction
(b) Avg. accuracy vs. number of model parameters (log scale).
@@
Source-Free Domain Adaptation (SFDA) aims to adapt a pre-trained model to a new target domain without requiring access to labeled source data. A common trend in SFDA is to use pre-trained models to bridge domain gaps and compensate for the absence of source data [SHOT_2021, kim2022source, ge2023domain, lai2023padclip, tang2023sourceMLLM]. Earlier works relied on ImageNet-pretrained backbones [SHOT_2021], while recent methods incorporate vision-language models (VLMs) like CLIP [singha2023ad, ge2023domain, lai2023padclip], using their transferable representations for pseudo-labeling [tang2023sourceMLLM], prompt tuning [lai2023padclip], or feature alignment [karim2023c]. However, these approaches typically focus on a single VLM, often involve prompt design or VLM training or finetuning.
While CLIP-based models offer strong visual-language alignment, their contrastive objectives limit semantic depth and reasoning ability. We ask a deeper question: can we move beyond shallow alignment to extract richer, instruction-following supervision from multimodal large language models (MLLMs)—models trained not just to align, but to understand, describe, and reason about images? To this end, we propose a new paradigm of foundation-model guided SFDA, where we treat MLLMs such as LLaVA [liu2023improvedllava] as frozen, zero-shot teachers. These models are trained on massive web-scale image-text corpora and encode diverse, high-level knowledge about visual concepts, object interactions, and real-world semantics, often combining high-capacity visual encoders with instruction-tuned LLMs, and recent studies show they now match or surpass CLIP on image classification benchmarks [liu2024revisiting]. Moreover, domain-specific MLLMs (e.g., medical [li2023llava-med]) make them a realistic choice when specialized teacher models are absent. As shown in Fig. 1(a), applying MLLMs in a zero-shot manner using our proposed STS (Sec. 3.1) module already outperforms SOTA SFDA methods. This validates their ability to generalize across domains without access to the source dataset and provides evidence that MLLMs encode transferable, domain-agnostic visual cues and reasoning that can guide adaptation.
However, directly using MLLMs in deployment is impractical. First, they are generative and not designed for closed-set classification, often violating instructions and producing free-form outputs. Second, as shown in Fig. 1(b), MLLMs are computationally expensive to query and unsuitable for real-time or on-device applications. Third, different MLLMs produce diverse and occasionally conflicting predictions, making it difficult to trust any single model’s output. This creates a dilemma: while MLLMs contain useful knowledge, their inference behavior is unreliable, and their scale makes them unsuitable for adaptation or downstream use.
To overcome these limitations, we introduce Reliability-based Curriculum Learning (RCL), a structured, multi-teacher SFDA framework that distills using pseudo-label guidance and self-learning from multiple frozen MLLMs into a compact student model. First, we use our proposed STS module to convert open-ended MLLM outputs into class predictions by computing semantic similarity between generated text and class names. Next, we estimate the reliability of pseudo-labels by measuring agreement across MLLMs, allowing us to partition unlabeled target samples based on consensus. RCL then trains the student model through a three-stage curriculum. In Reliable Knowledge Transfer (RKT), we use only highly consistent pseudo-labels to initialize adaptation. In Self-correcting and MLLM-guided Knowledge Expansion (SMKE), we incorporate partially ambiguous samples using model confidence and teacher agreement. Finally, Multi-hot Masking Refinement (MMR) addresses noisy supervision by masking low-confidence logits and applying consistency regularization. This staged design ensures stable convergence and maximizes utility from all target samples, while preserving scalability and efficiency. While conventional SFDA approaches rely on pretrained source models, our formulation follows recent trends [lai2023padclip, tang2023sourceMLLM] that treat foundation models as alternative supervisory signals. By distilling knowledge from multiple frozen MLLMs without requiring source data or finetuning, RCL remains faithful to the SFDA setting while enabling stronger supervision from multimodal sources.
RCL achieves SOTA performance on Office-Home, DomainNet-126, and VisDA-C, outperforming zero-shot MLLM baselines (with STS), majority-vote ensembles, and prior CLIP-based methods. RCL utilizes knowledge from MLLMs (70B to 7B params) into lightweight architectures like ResNet50 ( 25M params), achieving up to 280× to 2800× model size reduction while surpassing accuracy (see Fig. 1(b)). Our key contributions are as follows:
  • We propose the first SFDA framework that distills knowledge from multiple frozen MLLMs.
  • We introduce STS, a simple semantic matching method to convert open-ended MLLM outputs into class predictions.
  • We design RCL, a three-stage curriculum (RKT, SMKE, MMR) that incorporates pseudo-labels based on MLLM agreement and model confidence.
  • We achieve state-of-the-art results on Office-Home, DomainNet-126, and VisDA-C, outperforming CLIP-based and zero-shot MLLM baselines.
@@ 2 Related Work
Source-Free Domain Adaptation. SFDA adapts a pre-trained source model to a target domain using only unlabeled target data, making pseudo-labeling central [qu2022bmd, chenself, litrico2023guiding]. Prior works improve pseudo-labels via target structure [NRC_2022, tang2021nearest, tang2022semantic, diamant2024confusing, xu2025unraveling], align distributions [yang2021generalized, ding2022source], or generate target-style data [li2020model, VDMDA_arxiv_2021, 3CGAN_2022]. In contrast, several recent works adapt CLIP in DA/UDA settings with prompt learning, multi-prompt alignment [chen2023multi], prompt gradient alignment [hoang2024enhancing], prompt learning for DA [ge2023domain], and cluster-preserving prompt learning [vuong2025preserving]. These methods require source data or prompt optimization, whereas we tackle the stricter SFDA setting with single inference on teachers without any modification.
VLMs/MLLMs. Pretrained vision-language models (VLMs) like CLIP [CLIP_2021] and ALIGN [ALIGN_2021] capture vision-language-aligned features, while recent MLLMs (e.g., LLaVA [liu2023improvedllava], InstructBLIP [dai2024instructblip]) enhance multimodal understanding via LLM backbones, enabling strong zero-shot capabilities. While some works use CLIP for domain adaptation, they require labeled source data [lai2024empowering, zhao2024learning, westfechtel2023combining, hu2024reclip, zhang2025source] or fine-tuning [lai2023padclip]. DIFO [tang2023sourceMLLM], the latest SOTA work, adapts CLIP for target-domain learning by iteratively customizing it with prompt learning and distilling its knowledge into a task-specific model. While effective, this approach requires adapting a large VLM, whereas our method leverages multiple MLLMs in a zero-shot manner without finetuning.
Figure 2: Directly prompting MLLMs for classification can lead to failures: we propose semantic textual similarity (STS) in Section 3.1.
Multi-Teacher Knowledge Distillation. MTKD extends KD by combining knowledge from multiple teachers [jiang2024mtkd, pham2023collaborative], with strategies including aligning teacher outputs for consistent supervision [wu2021one], integrating teacher for better generalization [li2022mimicking], using heterogeneous teacher architectures [kang2023distillation], attention fusion [cheng2024lgfa], or dynamically balancing ensemble size via stochastic teacher selection [ding2024trade]. While these works highlight the benefits of combining multiple teachers, they generally assume labeled source data, stable teacher ensembles, and access to reliable outputs. RCL addresses the stricter SFDA setting, where no source data is available and teacher predictions (from MLLMs, CLIP, or others) can be inherently noisy. Our framework introduces consensus-based reliability scoring and a staged curriculum to progressively refine pseudo-labels, ensuring robust multi-teacher distillation under uncertainty.
3 Pseudo-labeling and Reliability
@@
3.1 Pseudo-labeling with MLLMs using STS
We use multiple MLLMs for initial target-image labeling. As MLLMs are designed primarily for text generation, their responses can deviate from classification. Therefore, we design prompts and reframe VQA as a zero-shot classification task using MLLMs (see Fig. 2). We design the prompt to incorporate all class names and a question instructing MLLMs to select the most appropriate match. The prompt and the image $x_{t}^{i}$ are then fed to multiple pre-trained MLLMs, such as LLaVA [liu2023improvedllava, liu2023llava], InstructBLIP [dai2024instructblip], and ShareGPT4V [chen2023sharegpt4v]. Each MLLM generates a text output $T_{1}^{i},T_{2}^{i},\ldots,T_{M}^{i}$, where $M$ is the number of MLLMs employed.
However, as shown in Fig. 2, MLLMs sometimes fail to follow prompts, generating responses beyond classification constraints. This happens when MLLMs rely on prior knowledge rather than selecting from provided options (e.g., predicting ‘Audi’ instead of ‘car’). Unlike MTKD, which distills structured soft labels, our approach must handle inherently diverse and inconsistent MLLM outputs. To mitigate this, we propose Semantic Textual Similarity (STS) to align outputs with class names, ensuring pseudo-label consistency. We derive pseudo-labels by computing STS between class names and MLLM-generated text. Formally, for the $m$-th MLLM and the $i$-th image $x_{t}^{i}$, the pseudo-label $\hat{y}^{mi}$ is determined by $\hat{y}^{mi}=\operatorname*{argmax}_{c}\text{STS}(T_{m}^{i},T_{t}^{c}),$ where $T_{t}^{c}$ represents the name of the $c$-th class. The STS between two text sequences $T_{1}$ and $T_{2}$ is computed as:
$\text{STS}(T_{1},T_{2})=\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{\|\mathbf{v}_{1}\|_{2}\|\mathbf{v}_{2}\|_{2}},$ (1)
where $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ are vector representations of $T_{1}$ and $T_{2}$, respectively. STS refines pseudo-labels by aligning MLLM outputs with class semantics, even when instructions are ignored. Implementation details and prompt sensitivity analysis are in Supplementary (Sec. 10).
Figure 4: Overview of our Reliability-based Curriculum Learning (RCL) framework. RCL applies curriculum learning over target data by progressively incorporating MLLM pseudo-labels based on reliability. (1) RKT uses high-confidence samples from full MLLM agreement to initialize feature learning. (2) SMKE integrates partially agreed labels to expand knowledge and correct early errors. (3) MMR learns from uncertain samples using multi-hot masking and consistency loss. This structured curriculum enables full data use while reducing label noise.
@@ 3.2 Consensus-based Reliability Measurement
Pseudo-labels from different MLLMs may vary for the same target sample. This disagreement raises a key question: how can we measure the reliability of the pseudo-labels from multiple MLLMs? While STS helps correct deviations from instructions, it does not assess pseudo-label reliability, as it cannot detect when MLLMs generate incorrect labels. To address this, we propose a consensus-based reliability metric for pseudo-labels. We define a reliability score $\mathcal{R}(x_{t}^{i})$ for each target domain sample $x_{t}^{i}$ based on agreement among MLLM-assigned pseudo-labels:
@@
where $\mathbbm{1}(\cdot)$ is the indicator function. $\mathcal{R}(x_{t}^{i})$ quantifies the proportion of MLLM pairs that assign the same pseudo-label to $x_{t}^{i}$. Using $\mathcal{R}(x_{t}^{i})$, we categorize $\mathcal{D}_{t}$ into three subsets: (1) Reliable ($\mathcal{D}_{\textit{R}}$): All MLLMs agree ($\mathcal{R}(x_{t}^{i})=1$), (2) Less Reliable ($\mathcal{D}_{\textit{LR}}$): Partial agreement ($0<\mathcal{R}(x_{t}^{i})<1$), and (3) Unreliable ($\mathcal{D}_{\textit{UR}}$): No agreement ($\mathcal{R}(x_{t}^{i})=0$). Thus, the target dataset is partitioned as: $\mathcal{D}_{t}=\{\mathcal{D}_{\textit{R}},\mathcal{D}_{\textit{LR}},\mathcal{D}_{\textit{UR}}\}.$
We observe that higher reliability scores correlate with improved pseudo-label accuracy, validating our consensus-based reliability metric. Fig. 3 further shows pseudo-label accuracy and sample distribution across reliability levels $\mathcal{R}(x_{t}^{i})$, highlighting disagreement among MLLMs. For samples with $\mathcal{R}(x_{t}^{i})>0$, accuracy is computed via majority vote across MLLMs, while for $\mathcal{R}(x_{t}^{i})=0$, we report the accuracy of each individual MLLM.